# BAnG: Bidirectional Anchored Generation for Conditional RNA Design

### v1

@@ Abstract
Designing RNA molecules that interact with specific proteins is a critical challenge in experimental and computational biology. Existing computational approaches require a substantial amount of experimentally determined RNA sequences for each specific protein or a detailed knowledge of RNA structure, restricting their utility in practice. To address this limitation, we develop RNA-BAnG, a deep learning-based model designed to generate RNA sequences for protein interactions without these requirements. Central to our approach is a novel generative method, Bidirectional Anchored Generation (BAnG), which leverages the observation that protein-binding RNA sequences often contain functional binding motifs embedded within broader sequence contexts. We first validate our method on generic synthetic tasks involving similar localized motifs to those appearing in RNAs, demonstrating its benefits over existing generative approaches. We then evaluate our model on biological sequences, showing its effectiveness for conditional RNA sequence design given a binding protein.
@@ 1 Introduction
The motivation for the design of the proposed generative method stems from two key observations. First, the total length of the RNA sequence to be generated is often unknown. Therefore, the method adopts an autoregressive approach. Second, RNA sequences that interact with proteins typically contain functional binding motifs —- specific regions that mediate interaction by forming molecular contacts with the protein. These binding motifs are embedded within larger sequence contexts, where the surrounding non-binding regions exert lesser influence on binding specificity. This makes it more effective to initiate sequence generation from the binding motif, rather than from the sequence’s ends, as is commonly done in current state-of-the-art NLP autoregressive models. These are the core ideas behind our method, Bidirectional ANchored Generation (BAnG).
@@ 2.1 Description of the BAnG generative approach
Figure 1: Schematic illustration of the RNA-BAnG generative process and its conditioning on the input protein 3D structure. The protein model was generated by AlphaFold2 and is colored by the secondary structure. RNA sequences are colored by nucleotides.
@@
In the BAnG framework, we leverage the following factorization of the joint distribution over a sequence $x=(x_{-n-1},...,x_{0},...,x_{n})$:
$P(x)$ $=P(x_{0})$
×$\quad\times\underbrace{\prod_{i=0}^{n}P(x_{-i-1}|x_{-i\dots i})}_{\text{left tokens}}\underbrace{\prod_{i=1}^{n}P(x_{i}|x_{-i\dots i-1})}_{\text{right tokens}},$ (1)
where $x_{0}$ is the first generated token. Concretely, sequence generation begins with two special anchor tokens, <ancl> and <ancr>, representing the left and right boundaries of the sequence, respectively. Tokens are then generated one at a time, alternating between directions: we first sample a token on the right, then a token on the left, and repeat this process to progressively extend the sequence outward. At each step, the distribution of the next token is conditioned on already generated ones, following (1). If an end-of-sequence (<eos>) token is generated in either direction, no further tokens are produced along that axis. The generation process stops when <eos> tokens are produced for both boundaries or when a predefined maximum sequence length is reached (see Figure 2).
@@
BAnG enables training a deep learning model to estimate the conditional distributions in (1) in order to perform bidirectional generation in the described manner, using a process analogous to autoregressive training. The key difference in the architecture compared to the standard autoregressive case is the replacement of the conventional lower triangular attention mask with a specifically designed bidirectional attention mask, shown in Figure 3. This custom mask ensures that any representation of a given token cannot depend on tokens beyond those appearing in the corresponding conditional in (1). This simple modification allows the model to learn the dependencies needed for our generation strategy, while ensuring efficient parallelization of the forward and backward passes across all tokens in a sequence during training. During model training and inference, the probabilities for the next token in each direction are derived from the embedding of the most recently generated token in the same direction, as shown schematically in Figure 4.
@@
We shall emphasize that the single pass training with the conditional factorization in (1) is only possible thanks to the introduction of two anchor tokens. Indeed, if only a single anchor token was used, it would be responsible for predicting the first tokens in both directions, right and left. To prevent information leakage, its attention would be restricted to itself, which would make the prediction of the left token independent of the right one. This lack of conditioning could lead to the generation of incompatible token pairs.
@@ 2.2 Model
RNA-BAnG architecture consists of two main components: a protein module and a nucleotide module. The protein module derives a representation from the protein’s sequence and structure, while the nucleotide module generates a nucleotide sequence conditioned on this representation.
@@
Our model’s modules comprise several main blocks - Embedder, Self Attention, Geometric Attention, and Cross Attention, schematically illustrated in Figure 5. The Embedder block generates token embeddings from protein, RNA, and DNA sequences. The inclusion of RNA and DNA sequences in the training data serves to augment the dataset, as it can help the model learn the shared patterns. The Embedder block processes amino acids with one-hot encoding residue types (20 canonical and a padding token), followed by a linear transformation and normalization of the resulting embeddings. For nucleic acids, it encodes sequence information in two steps. First, it one-hot encodes residue types, using tokens for the four standard RNA nucleotides, along with two anchor tokens, <eos> token, and a padding (<pad>) token. DNA residues are treated as RNA equivalents, with thymine (T) replaced by uracil (U). Second, the sequence type (DNA or RNA) is one-hot encoded, linearly transformed, and concatenated with the residue embeddings. This combined representation is then passed through a normalization layer.
@@
Geometric Attention aims to incorporate protein structural information. To achieve this, we adopted the protein representation introduced in AlphaFold2 (Jumper et al., 2021). Specifically, for each protein residue, a rigid frame $T$ is constructed based on the coordinates of its $C$, $C_{\alpha}$, and $N$ backbone atoms. Attention between these frames is then calculated using the geometric part from AlphaFold2’s Invariant Point Attention (IPA) mechanism (more details in Appendix A). The importance of this module is demonstrated by the fact that RNA-BAnG fails to converge in its absence.
@@ 3 Validating BAnG on a synthetic task
To evaluate the BAnG method, we first consider a synthetic task that emulates a real-world conditional generation scenario. As mentioned above, RNA sequences often include a functional binding motif within a broader sequence context, where non-binding regions contribute minimally to binding specificity (Ray et al., 2013). Based on this observation, the task’s objective is to generate sequences that contain a predefined short subsequence, synthetic motif.
@@
Reference methods.
Evaluation results.
With each tested approach we generated 1,000 sequences. Table 1 summarizes the performance of each method. Additional statistics on the frequency of each synthetic motif in the generated data for DoubleBind task are provided in Table 2. Examples of generated sequences can be found in Section B.3.
@@
The tables show that BAnG outperforms other methods, with a particularly notable margin on the DoubleBind task. BAnG also generates fewer sequences containing mixed synthetic motifs. The very low performance of ESM3’s iterative methods is expected, as the lack of absolute positional dependencies in the data causes the model to assign nearly uniform probabilities across tokens. This reasoning is further supported by the significant performance improvement observed when positional information is introduced through the anchor token in IAnG. Nonetheless, both methods suffer from a key limitation: a discrepancy between the demasking process during training and inference, which undermines their overall effectiveness.
@@ 4.1 Data
We collected our protein-nucleotide interaction data from the Protein Data Bank (PDB) (Berman, 2000), utilizing information provided in the PPI3D database (Dapkūnas et al., 2024). However, we conducted independent postprocessing of the data, distinct from PPI3D, as they focus on structural RNA and DNA information, while we are concerned solely with their sequences. The postprocessing steps involved verifying chain interactions, discarding ambiguous protein structures, and excluding chains containing non-standard residues, as explained in more detail in Section C.1. During training and validation, each time a sample was encountered, its anchor point was randomly sampled from the interacting nucleotides. The anchor tokens were then inserted right after the selected nucleotide. Additionally, to diversify RNA sequence information, we collected non-coding sequences from RNAcentral (release 24), a comprehensive database integrating RNA sequences from multiple expert sources (The RNAcentral Consortium et al., 2019) (more details in Section C.2). Since we lack information about their interactions, the anchor points for solo RNA sequences were selected randomly.
@@ 4.2 RNA-BAnG training
We designed the training process to consist of two steps: the first for the model to learn general information about RNA sequences, and the second for the model to learn conditioning on proteins. In both steps, the training objective is the cross-entropy loss between the predicted and ground truth nucleotide token probabilities (Bengio et al., 2000). As the first step we train RNA-BAnG nucleotide module without cross-attention block on standalone RNA sequence data from RNAcentral. During training, the loss values for the four tokens closest to each anchor on either side are weighted at 0.01. This weighting scheme is applied because the model lacks the context needed to accurately predict these residues. Next, we train the full model, using weights from the previous step, on the combined protein-nucleotide sequence data, this time without applying any loss weighting. Additional training details can be found in Appendix D.
@@ 4.3 Evaluation protocol
To evaluate the performance of our model, we adopted the scoring approach proposed by the authors of GenerRNA (Zhao et al., 2024) and others (Im et al., 2019). This method leverages DeepCLIP (Grønning et al., 2020), a state-of-the-art predictive model that, after being trained on examples of interacting and non-interacting RNA sequences for a given protein, can assign binding probabilities to any RNA sequence with the same protein. These probabilities serve as a proxy for evaluating the quality of sequences generated by RNA-BAnG and other methods.
To obtain suitable training and testing data for DeepCLIP, we utilized RNAcompete experiments (Ray et al., 2009), conducted by the authors of the RNA Compendium (Ray et al., 2013). Using this data, we trained DeepCLIP models for each sample and selected only those that met our performance criteria (see Section E.1). Separately, for each sample, we derived 1,000 sequences identified by RNAcompete as highly likely to bind to the protein (positive examples) and another 1,000 sequences as highly unlikely to bind (negative examples), based on their experimental affinity scores (see Section E.1). We excluded these sequences from the DeepCLIP model training, ensuring that they served as an independent benchmark for testing the quality of both the DeepCLIP predictions and the RNA sequences generated by RNA-BAnG. Below, we refer to them as to the positive and negative experimental sets.
Due to the lack of experimentally solved protein structures for the test set, and aiming for a more robust method, we only used AlphaFold2 protein models for RNA-BAnG inference during evaluation. Concretely, for each RNA Compendium sample, we generated the corresponding three-dimensional protein model, and retained for further testing only those with a predicted local distance difference test (pLDDT) score greater than 70%. This threshold is well-accepted in the community and guarantees structural reliability of the produced protein structures, at least at the domain level. The final test set consisted of 71 samples, representing 67 unique protein sequences.
To match the amount of RNA sequences in experimental sets, we generated 1,000 of them with RNA-BAnG for each test sample. We also generated 1,000 random sequences to serve as the simplest baseline (see Section E.2).
@@ 4.4 Experiment results
Averaging across samples conceals important details about the statistics of the compared sequences. For a more detailed, sample-by-sample analysis, shown in Figure 7, we selected a threshold value of 0.75, as DeepCLIP scores around 0.5 can be ambiguous in interpretation. We consider sequences with DeepCLIP scores above this value as high-affinity RNA sequences (haRNA). As one can see from Figure 7, RNA-BAnG generates more high-affinity sequences than a random generator for 56 out of 71 test samples. For 33 of the test samples, more than half of the sequences generated by RNA-BAnG are the high-affinity ones. Additionally, our model generated less than 20% high-affinity sequences for only 14 out of 71 test samples. The high success rate of randomly generated sequences for some samples may be attributed to the low complexity of captured protein binding motifs (Ray et al., 2013). The examples of generated sequences for some representative samples and individual DeepCLIP score distributions can be found in Appendix F.
The performance of our model shows no significant correlation with the predicted quality of the protein 3D model (pLDDT) or with the sequence similarity of a protein target to the training data (Figure F.1 in Appendix F). This analysis suggests that the model does not only memorize the information, but exhibits a degree of generalization. We can also conclude that the performance accuracy of RNA-BAnG can not be attributed to protein structure prediction quality, if its pLDDT score is higher than 70%.
@@
The low performance of RNAFLOW can be explained by a discrepancy between its training process and the inference we conducted. Concretely, RNAFLOW was trained on protein sequences and structures truncated to 50 residues, ensuring the inclusion of the protein binding site. However, we tested it on proteins of varying lengths, ranging from 177 to 589 amino acids, without highlighting their binding sites through truncation (since this information is unavailable). Additionally, the proteins we tested had zero sequence similarity to those in RNAFLOW’s training set (PDB), making it more challenging to predict their structure in complex with RNA. These limitations significantly restrict the applicability of RNAFLOW to RNA sequence design, especially when compared to RNA-BAnG.

### v2

@@ Abstract
Designing RNA molecules that interact with specific proteins is a critical challenge in experimental and computational biology. Existing computational approaches require a substantial amount of previously known interacting RNA sequences for each specific protein or a detailed knowledge of RNA structure, restricting their utility in practice. To address this limitation, we develop RNA-BAnG, a deep learning-based model designed to generate RNA sequences for protein interactions without these requirements. Central to our approach is a novel generative method, Bidirectional Anchored Generation (BAnG), which leverages the observation that protein-binding RNA sequences often contain functional binding motifs embedded within broader sequence contexts. We first validate our method on generic synthetic tasks involving similar localized motifs to those appearing in RNAs, demonstrating its benefits over existing generative approaches. We then evaluate our model on biological sequences, showing its effectiveness for conditional RNA sequence design given a binding protein.
@@ 1 Introduction
The motivation for the design of the proposed generative method stems from two key observations. First, the total length of the RNA sequence to be generated is often unknown. Therefore, the method adopts an autoregressive approach. Second, RNA sequences that interact with proteins typically contain functional binding motifs – specific regions that mediate interaction by forming molecular contacts with the protein. These binding motifs are embedded within larger sequence contexts, where the surrounding non-binding regions exert lesser influence on binding specificity (Ray et al., 2013). This makes it more effective to initiate sequence generation from the binding motif, rather than from the sequence’s ends, as is commonly done in current state-of-the-art NLP autoregressive models. These are the core ideas behind our method, Bidirectional ANchored Generation (BAnG).
Figure 1: Schematic illustration of the RNA-BAnG generative process and its conditioning on the input protein 3D structure. The protein model was generated by AlphaFold2. RNA sequences are colored by nucleotides.
@@ 2.1 Description of the BAnG Generative Approach
Our approach models conditional distributions of each token given previously generated ones, similar to autoregressive modeling, but with a different factorization of the joint distribution over the sequence, which leads to a different order of generation. In the BAnG framework, we leverage the following factorization of the joint distribution over a sequence $x=(x_{-m},...,x_{0},...,x_{n})$:
$P(x)=P(x_{0})\underbrace{\prod_{i=1}^{m}P(x_{-i}|\mathcal{C}^{l}_{-i})}_{\text{left tokens}}\underbrace{\prod_{i=1}^{n}P(x_{i}|\mathcal{C}^{r}_{i})}_{\text{right tokens}},$ (1)
where $\mathcal{C}^{l}_{-i}=(x_{-i+1},\dots,x_{i})$ and $\mathcal{C}^{r}_{i}=(x_{-i+1},\dots,x_{i-1})$ denote left and right contexts. This formulation enables generation to proceed outward from the token $x_{0}$, conditioning each new token on a growing window of previously generated tokens. While the factorization is centered on $x_{0}$, we still need some initial tokens to begin generation. To this end, we insert two special anchor tokens, <ancl> and <ancr>, immediately before $x_{0}$: $x=(x_{-m},...,\texttt{<ancl>},\texttt{<ancr>},x_{0},...,x_{n})$. These tokens provide a fixed starting context, making it possible to assign a probability to $x_{0}$ and begin outward generation. They remain in the contexts $\mathcal{C}^{l}_{-i}$ and $\mathcal{C}^{r}_{i}$ for all subsequent tokens.
With this setup in place, generation proceeds token by token, alternating between the right and left directions: we first sample a token on the right, then a token on the left, and repeat this process to progressively extend the sequence outward. At each step, the distribution of the next token is conditioned on already generated ones, following Eq. 1. If an end-of-sequence (<eos>) token is generated in either direction, no further tokens are produced along that axis. The generation process stops when <eos> tokens are produced for both boundaries or when a predefined maximum sequence length is reached (see Figure 2).
@@
BAnG enables training a deep learning model to estimate the conditional distributions in Eq. 1 in order to perform bidirectional generation in the described manner, using a process analogous to autoregressive training. The key difference in the architecture compared to the standard autoregressive case is the replacement of the conventional lower triangular attention mask with a specifically designed bidirectional attention mask, shown in Figure 3. This custom mask ensures that any representation of a given token cannot depend on tokens beyond those appearing in the corresponding conditional in Eq. 1. This simple modification allows the model to learn the dependencies needed for our generation strategy, while ensuring efficient parallelization of the forward and backward passes across all tokens in a sequence during training. During model training and inference, the probabilities for the next token in each direction are derived from the embedding of the most recently generated token in the same direction, as shown schematically in Figure 4.
@@
We shall emphasize that the single pass training with the conditional factorization in Eq. 1 is only possible thanks to the introduction of two anchor tokens. Indeed, if only a single anchor token was used, it would be responsible for predicting the first tokens in both directions, right and left. To prevent information leakage, its attention would be restricted to itself, which would make the prediction of the left token independent of the right one. This lack of conditioning could lead to the generation of incompatible token pairs.
@@ 2.2 Model
The RNA-BAnG architecture consists of two main components: a protein module and a nucleotide module. The protein module derives a representation from the protein’s sequence and structure, while the nucleotide module generates a nucleotide sequence conditioned on this representation. Our model’s modules comprise several main blocks - Embedder, Self Attention, Geometric Attention, and Cross Attention, schematically illustrated in Figure 5. The Embedder block generates token embeddings from protein, RNA, and DNA sequences. The inclusion of RNA and DNA sequences in the training data serves to augment the dataset, as it can help the model learn the shared patterns. Protein sequences are tokenized using the 20 standard amino acids and a padding token. RNA sequences are tokenized using the four canonical nucleotides, along with padding and end-of-sequence tokens. DNA sequences are tokenized with the same set, treating DNA residues as their RNA equivalents. To allow the model to distinguish between RNA and DNA, additional sequence type information is encoded separately (more details in Appendix A).
@@
Geometric Attention aims to incorporate protein structural information. To achieve this, we adopted the protein representation introduced in AlphaFold2 (Jumper et al., 2021). Specifically, for each protein residue, a rigid frame $T$ is constructed based on the coordinates of its $C$, $C_{\alpha}$, and $N$ backbone atoms. Attention between these frames is then calculated using the geometric part from AlphaFold2’s Invariant Point Attention (IPA) mechanism (more details in Appendix A).
@@ 3 Validating BAnG on a Synthetic Task
To evaluate the BAnG method, we first consider a synthetic task that emulates a real-world conditional generation scenario. As mentioned above, RNA sequences often include a functional binding motif within a broader sequence context, where non-binding regions contribute minimally to binding specificity. The binding motifs are typically short and linear, with some being highly conserved and others exhibiting probabilistic patterns (Ray et al., 2013). Based on this observation, the task’s objective is set to generate sequences that contain a predefined short subsequence, synthetic motif. Although these simplifications do not fully capture the nuances of real-world scenarios, they offer a baseline for understanding the method before applying it to more challenging tasks.
@@
Reference methods.
@@
Evaluation results.
With each tested approach, we generated 1,000 sequences. Table 1 summarizes the performance of each method. Additional statistics on the frequency of each synthetic motif in the generated data for DoubleBind task are listed in Table 2. Examples of generated sequences can be found in Section B.3.
The tables show that BAnG outperforms other methods, with a particularly notable margin on the DoubleBind task. BAnG also generates fewer sequences containing mixed synthetic motifs. We have also conducted experiments demonstrating BAnG’s robustness to both larger fixed sequence lengths and varying sequence lengths, showing consistent performance across different configurations (Section B.4).
The very low performance of ESM3’s iterative methods is expected, as the lack of absolute positional dependencies in the data causes the model to assign nearly uniform probabilities across tokens. This reasoning is further supported by the significant performance improvement observed when positional information is introduced through the anchor token in IAnG. Nonetheless, both methods suffer from a key limitation: a discrepancy between the demasking process during training and inference, which undermines their overall effectiveness.
@@ 4.1 Data
We collected our protein-nucleotide interaction data from the Protein Data Bank (PDB) (Berman, 2000), utilizing information provided in the PPI3D database (Dapkūnas et al., 2024). However, we conducted independent postprocessing of the data, distinct from PPI3D, as they focus on structural RNA and DNA information, while we are concerned solely with their sequences. The postprocessing steps involved verifying chain interactions, discarding ambiguous protein structures, and excluding chains containing non-standard residues, as explained in more detail in Section C.1. During training and validation, each time a sample was encountered, its anchor point was randomly sampled from the interacting nucleotides. The anchor tokens were then inserted right after the selected nucleotide. Additionally, to diversify RNA sequence information, we collected non-coding sequences from RNAcentral (release 24), a comprehensive database integrating RNA sequences from multiple expert sources (The RNAcentral Consortium et al., 2019) (more details in Section C.2). Since interaction data for these sequences is unavailable, their anchor points were selected randomly.
@@ 4.2 RNA-BAnG Training
We designed the training process to consist of two steps: the first for the model to learn general information about RNA sequences, and the second for the model to learn conditioning on proteins. In both steps, the training objective is the cross-entropy loss between the predicted and ground-truth nucleotide token probabilities (Bengio et al., 2000). As the first step we train RNA-BAnG nucleotide module without cross-attention block on standalone RNA sequence data from RNAcentral. During training, the loss values for the four tokens closest to each anchor on either side are weighted at 0.01. This weighting scheme is applied because the model lacks the context needed to accurately predict these residues. Next, we train the full model, using weights from the previous step, on the combined protein-nucleotide sequence data, this time without applying any loss weighting. Additional training details can be found in Appendix D.
@@ 4.3 Evaluation Protocol
Evaluating generative models is difficult because direct comparison with ground truth is not possible. To the best of our knowledge, there are no established computational protocols for measuring RNA affinity to proteins. Therefore, we use a predictive model as our main evaluation tool, following common practice in machine learning, such as FID scores in image generation or refoldability metrics in protein design.
Scoring model.
As the scoring model, we adopted DeepCLIP (Grønning et al., 2020), following the approach of the authors of GenerRNA (Zhao et al., 2024) and others (Im et al., 2019). DeepCLIP is a state-of-the-art predictive model that, after being trained on examples of interacting and non-interacting RNA sequences for a given protein, can assign binding probabilities with the same protein to any RNA sequence. These probabilities serve as a proxy for evaluating the quality of sequences generated by RNA-BAnG and other methods.
Test set.
We built our test set using data from RNAcompete experiments (Ray et al., 2009), conducted by the authors of the RNA Compendium (Ray et al., 2013). Each sample consists of a protein sequence paired with a set of RNA sequences, each assigned a score reflecting its binding affinity to the protein. We trained a separate DeepCLIP model for each sample and excluded those where the scoring model did not meet our performance criteria (see Section E.1).
Due to the lack of experimentally solved protein structures for the test set and in pursuit of a more robust method, we used only AlphaFold2 protein models for RNA-BAnG inference during evaluation. For each sample in our test set, we generated the corresponding three-dimensional protein model and retained only those with a predicted local distance difference test (pLDDT) score greater than 70%. This threshold, widely accepted in the community, ensures the structural reliability of the produced protein models, at least at the domain level. The final test set included 71 samples, representing 67 unique protein sequences.
To evaluate the generative models in a relative way, we built positive and negative sets of RNA sequences for each sample. The positive set includes sequences with the highest experimental affinity scores, while the negative set includes those with the lowest. We excluded these sets from DeepCLIP training and used them as an independent benchmark to assess both DeepCLIP predictions and the quality of the generated RNA sequences (see Section E.1 for details). In addition to generating sequences with the evaluated model, we also created random sequences to serve as a baseline (see Section E.2). Each mentioned set contains 1,000 sequences, unless stated otherwise.
@@ 4.4 Experiment Results
To complement the threshold-dependent performance analysis and capture variability concealed by averaging across samples, we conducted a sample-by-sample evaluation. Using a fixed threshold, we classified sequences as high-affinity RNA (haRNA) if their assigned score exceeded this value. For each sample in the test set, we then quantified the proportion of sequences meeting the haRNA criterion, presented in Figure 7. The fixed threshold value of 0.75 was chosen as detailed in the Appendix F. As one can see from Figure 7, RNA-BAnG generates more high-affinity sequences than a random generator for 56 out of 71 test samples. For 33 of the test samples, more than half of the sequences generated by RNA-BAnG are the high-affinity ones. Additionally, our model generated less than 20% high-affinity sequences for only 14 out of 71 test samples. The high success rate of randomly generated sequences for some samples may be attributed to the low complexity of captured protein binding motifs (Ray et al., 2013).
The performance of our model shows no significant correlation with the predicted quality of the protein 3D model (pLDDT) or with the sequence similarity of a protein target to the training data (Figure F.2 in Appendix F). This analysis suggests that the model does not only memorize the information, but exhibits a degree of generalization. We can also conclude that RNA-BAnG’s performance is robust to the quality of protein structure prediction, provided the pLDDT score exceeds 70%.
In addition to the deep learning-based scoring model, we explored several computational approaches to evaluate the generated sequences. Among these, one method yielded results that showed a meaningful correlation with the deep learning scores, supporting the validity of the model’s output. Detailed descriptions of the computational analyses, along with examples of generated sequences for representative samples and individual DeepCLIP score distributions, are provided in Appendix F.
@@
The low performance of RNAFLOW can be explained by a discrepancy between its training process and the inference we conducted. Concretely, RNAFLOW was trained on protein sequences and structures truncated to 50 residues, ensuring the inclusion of the protein binding site. However, we tested it on proteins of varying lengths, ranging from 177 to 589 amino acids, without highlighting their binding sites through truncation (since this information is unavailable). Moreover, the proteins we tested had zero sequence similarity to those used to train RNA-BAnG. As both RNA-BAnG and RNAFLOW are trained on PDB-derived data, it is reasonable to assume low similarity to RNAFLOW’s training set as well. This could make prediction of their structure in complex with RNA more challenging. These limitations significantly restrict the applicability of RNAFLOW to RNA sequence design, especially when compared to RNA-BAnG.